OpenStudy (anonymous):

For example if a test has 22 questions and was worth 173 points and you got 30 points wrong what would your grade be?

OpenStudy (superdavesuper):

Need more info - how many grades are there for starter?

OpenStudy (anonymous):

what do you mean cause im looking to see if the grade would be an A, B, C, D, or F

OpenStudy (nikato):

173-30 --------- 173

OpenStudy (superdavesuper):

173/5=34.6....so Grade F would be 0-34, D 35-69 and so on? @nikato hahaa yup that is the right calculation :)

OpenStudy (anonymous):

They grading is by A = 90-100% B = 80-89% C = 70-79% D = 60-69% F = 59% and below

OpenStudy (nikato):

So did u solve that fraction? Then change it to a decimal

OpenStudy (anonymous):

you got 143 out of 173 which mean you got about 18 question correct out of 22! so your grade is a B-

OpenStudy (anonymous):

Thanks what I wanted to know

OpenStudy (anonymous):

aha bro, anytime :)

OpenStudy (anonymous):

but how did you get that @LaiRamJ

OpenStudy (anonymous):

like what is the math on how to get that answer

OpenStudy (anonymous):